calculate Vr
How would you estimste the error in your experimental voltage and experimental time values?
4000F 3120 4.105 V capacilor, 400 F 1 = 83.14 ms tirne step = 5 jas RC = 48 ms
Answer #1

V = V-ve =v-[vci-e te t v=v-vtve Rc ند e ve the R t = 41105 e 48ms = 4.105.e looot 48 x = 20.834 - 4.105 efor error calculation data on graph required